OpenStudy (anonymous):

What is the ratio 60/135 written in lowest terms?

jigglypuff314 (jigglypuff314):

Hello and Welcome to Openstudy! :) well 60 and 135 are both dividable by 5 right? so 60 divided by 5 = 12 and 135 divided by 5 = 27 so you get 12/27 then 12 and 27 are then both dividable by 3 right? so 12 divided by 3 = 4 and 27 divided by 3 = 9

OpenStudy (triciaal):

write the prime factors of numerator/prime factors of denominator for this problem 60/135 60 = 2.2.3.5 135 =3.3.3.5 cancel what is common in both (3.5) 2.2/3.3 = 4/9